OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

apply two point form fomula............for this .......

OpenStudy (mertsj):

Find the slope. \[m=\frac{8-0}{2-(-2)}=\frac{8}{4}=2\]

OpenStudy (mertsj):

Now choose one of the two points. I choose (-2,8) and plug into point-slope form: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (mertsj):

\[y-8=2(x-(-2))\] \[y-8=2(x+2)\] \[y-8=2x+4\] \[y=2x+12\]